OpenStudy (anonymous):

why does sin(-.001) show up as -.001 on my ti-nspire cx calculator?

OpenStudy (anonymous):

because if \(x\) is very close to \(0\) then \(\sin(x)\) is very close to \(x\)

OpenStudy (anonymous):

which is another way of saying \[\lim_{x\to 0}\frac{\sin(x)}{x}=1\]